OpenStudy (anonymous):

PLEASE HELP Let n be the middle number of three consecutive even integers. Write an expression for the sum of these integers.

OpenStudy (anonymous):

even numbers are per definition, divisible by 2 right ?

OpenStudy (anonymous):

the equation is 3x+6 where x is the first number

OpenStudy (anonymous):

says be sure to use a variable

OpenStudy (anonymous):

a correct* variable

OpenStudy (anonymous):

wait i got the answeer ...

OpenStudy (anonymous):

the answer is 3n

OpenStudy (anonymous):

if n is the middle term then the term greater to it will be n+2 and the term lesser to it will be n-2 ... when you add then all you get n-2+n+n+2=3n

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah,that's it. the knowledge is that even numbers count +2,+2,+2: 2,4,6,8,... number * 2 will be divisible by 2, and one more *2 is +2

OpenStudy (anonymous):

3*2 = 6, 4*2 = 8 - which is the same as 6+2

OpenStudy (anonymous):

i understand it now ty